# Contact Wasserstein Geodesics for Non-Conservative Schrödinger Bridges

### v2

@@ 1 Introduction
Inferring the stochastic process that most likely generates a set of sparse observations is a fundamental challenge, e.g., in cellular dynamics (yeo2021generative; zhang2024scnode; moon2019visualizing), meteorology (franzke2015stochastic), and economics (kazakevivcius2021understanding; huang2024generative). Here, the target is not merely the distributions of observed data, but rather the underlying dynamics of cell populations, weather patterns, or economic phenomena, enabling reconstruction of missing intermediate states and predicting the systems’ future evolution.
The Schrödinger Bridge (SB, schrodinger1931umkehrung) is a powerful mathematical framework to address this. SB seeks the most likely stochastic path between marginals (i.e., observations), while being close to a reference process, typically Brownian motion. This offers a general stochastic optimal control perspective that encompasses both Optimal Transport (OT, vargas2021solving) and generative approaches such as diffusion models (ho2020denoising; chen2024overview), which can be interpreted as optimal bridges with a Gaussian initial marginal (bortoli2021diffusion). Unfortunately, current SB solvers operate on an infinite probability space and rely on iterative forward–backward stochastic simulations or progressive refinement of the reference dynamics. This leads to complex and costly optimizations, limiting adoption.
@@
This paper provides a novel mathematical generalization of the SB to model non-conservative systems and develops near-linear time algorithms to realize the framework. We build on the geometric perspective of the SB, which casts it as a flow governed by Hamiltonian dynamics on the Wasserstein probability space (Sec. 3). Extending the Hamiltonian system to a contact Hamiltonian (zadrathesis), we propose a more general energy-varying formulation of the SB problem: The Non-Conservative Generalized SB (NCGSB, Sec. 4). To make computations tractable, we introduce the Contact Wasserstein Geodesic (CWG), which solves the NCGSB problem by casting it as a geodesic computation, reshaping its cost functional into a Riemannian metric whose induced distance is minimized. Discretizing the geodesic leads to geodesic segments that match standard residual blocks. We show how this leads to an efficient solver that avoids outer iteration loops and achieves near-linear complexity in both dimensionality and batch size. Additionally, our approach allows for metric modulation, enabling guided generation tailored to the task’s specifications (Sec. 5). We demonstrate our approach on benchmarks and tasks such as LiDAR manifold navigation, molecular dynamics predictions, and image-based reconstruction of systems such as sea-surface temperature and robotic pick-and-place (Sec. 6).
In summary, we contribute: (1) a novel non-conservative formulation of the Schrödinger Bridge problem that models a wider range of real-world physical stochastic processes; (2) the introduction of the Contact Wasserstein Geodesic (CWG) framework, a general geometric solver compatible with all Schrödinger Bridge variants, enabling efficient and scalable computation; (3) a guided generation methodology based on the modulation of the metric associated with CWG.
@@ 2 Related Work
Schrödinger bridges.   The SB problem imposes no constraints on the probabilistic path beyond matching the endpoint marginals, limiting its applicability when intermediate observations exist. It also does not permit including known physical laws governing the system’s dynamics. The multi-marginal Schrödinger Bridge (mmSB) treats intermediate observations (theodoropoulos2025momentum) as constraints, which enables reconstruction of continuous dynamics without piecewise approximations. In contrast, the Generalized Schrödinger Bridge (GSB) adds a state cost, allowing potential energy functionals to be minimized along the probability path. This let us to model mean-field interactions (gaitonde2021polarization; ruthotto2020machine), conservative forces (philippidis1979quantum; noe2020machine), or geometric priors (chen2024flow; liu2018rob).
@@
Non-conservative Schrödinger bridge formulations.   The SB problem assumes constant energy, preventing it from modeling non-conservative systems. The momentum SB augments the state space with velocity (theodoropoulos2025momentum; chen2023deep), allowing damping to be incorporated (blessing2025underdamped; sterling2025critically). This, however, doubles the state space and increases computational cost. Other extensions replace the Brownian reference process with the Ornstein–Uhlenbeck (OU) process (orland2025schrodinger; zhang2025learning), introducing a non-conservative prior for the target dynamics. The OU process defines a mean distribution with a curl-free component that drives convergence and a divergence-free component that induces rotation. Yet, this formulation lacks a mechanism for energy dissipation in the rotational dynamics. We introduce a more general energy-varying framework in which dissipation naturally emerges across all components of the dynamics, while only requiring a scalar augmentation of the state space.
@@
Schrödinger bridge solvers.   Matching-based iterative approaches (shi2023diffusion; gushchin2024light; peluchetti2023diffusion) have gained popularity by improving the scalability and robustness of traditional Iterative Proportional Fitting (IPF) algorithms (kullback1968probability; léonard2013surveyschrodingerproblemconnections) through Markovian projections, thereby avoiding the need for full trajectory storage. However, for GSB, restrictive assumptions like Gaussian probability paths limit expressivity (liu2024generalized; tang2025branched), while for mmSB, global trajectory consistency is compromised due to the piecewise nature of the approach and its sensitivity to the initial choice of the reference process (shen2025multimarginal). To overcome these limitations, the stochastic dynamics can be learned indirectly by leveraging the analytical optimality conditions of the SB problem (vargas2021solving; chen2022likelihood). This has been shown to scale more effectively to both the GSB (liu2022deep; buzun2025hotahamiltonianframeworkoptimal) and the mmSB problem (theodoropoulos2025momentum; hong2025trajectory; chen2023deep), due to the symmetries inherent in these optimality conditions (App. A). However, these methods are limited by a classical SB solver bottleneck: the computational overhead of their iterative nature, which alternates between forward–backward passes or repeated dynamics integration and reference updates. We propose a cheaper non-iterative solver that scales nearly linearly with both dimensionality and sample size.
@@
Schrödinger bridge guided generation.   The SB framework naturally extends to conditional generation, where the marginals and the transition path depend on additional parameters or objectives (shi2022conditional). Model guidance techniques (pmlr-v202-song23k; guo2024gradient) introduce a guidance term into the stochastic process, typically derived from the gradient of a loss function. This approach steers the flow locally and resembles a gradient-based form of optimal control. Alternatively, raja2025actionminimization employs a global optimal control perspective. Their approach generates full trajectories and chooses the one minimizing a task-specific action functional. However, their method is deterministic and yields a single optimal path rather than a posterior distribution over paths. Unlike previous guidance methods, we propose a hybrid approach to guided generation within the NCGSB framework. By embedding a task-specific loss into the potential, we reshape the Riemannian metric so that the resulting geodesic reflects the guidance objective, allowing the learned dynamics to align with the desired outcome.
@@ 3 Preliminaries
The Wasserstein manifold.   Before formally introducing the SB problem, we define its domain. Let $\mathcal{P}^{+}(\mathcal{M})$ denote the space of smooth, positive density functions supported on a manifold $\mathcal{M}$. Each element $\rho\in\mathcal{P}^{+}(\mathcal{M})$ is a function $\rho(x):\mathcal{M}\to\mathbb{R}^{+}$ satisfying $\int_{\mathcal{M}}\rho(x)\,dx=1$. The density dynamics is represented by a time-dependent family $\{\rho^{t}\}_{t\in\mathbb{R}^{+}}\subset\mathcal{P}^{+}(\mathcal{M})$. The infinitesimal variation of the density at time $t$ is the time derivative $\partial_{t}\rho^{t}(x)$, which lies in the tangent space $\mathcal{T}_{\rho^{t}}\mathcal{P}^{+}(\mathcal{M})$. The collection of all such tangent spaces forms the tangent bundle $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$. When equipped with the Wasserstein metric, $\mathcal{P}^{+}(\mathcal{M})$ becomes a Riemannian manifold (ambrosio2005gradient). The corresponding metric tensor is defined as,
@@
where $(-\Delta_{\rho^{t}})^{\dagger}$ is the inverse of the weighted Laplacian operator $\Delta_{\rho^{t}}=-\nabla_{x}\cdot(\rho^{t}\,\nabla_{x})$ (chow2020wasserstein), inducing an inner product on $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ and a distance $d_{\mathcal{W}_{2}}(\rho_{a},\rho_{b})$ for $\rho_{a},\rho_{b}\in\mathcal{P}^{+}(\mathcal{M})$. The minimum-length curve $\rho^{t}$ connecting the two distributions $\rho_{a},\rho_{b}$ is called a geodesic.
Multi-marginal generalized Schrödinger bridge (mmGSB).   Given two endpoint densities $\rho_{a},\rho_{b}\in\mathcal{P}^{+}(\mathcal{M})$, the SB problem (schrodinger1931umkehrung; schrodinger1932theorie) seeks the most probable interpolating density path $\rho^{t}$. This minimizes the Kullback–Leibler divergence w.r.t. a reference process $\rho^{t}_{\text{ref}}$, typically Brownian motion. The SB problem is equivalent to a stochastic optimal control setting (dai1991stochastic), which minimizes the cost required to transport a set of diffusing particles from an initial distribution $\rho_{a}$ to a target distribution $\rho_{b}$. This dynamic reformulation of the OT problem (benamou2000computational; chen2014relationoptimaltransportschrodinger) has solutions corresponding to geodesics on the Wasserstein manifold $\mathcal{P}^{+}(\mathcal{M})$. These trajectories are straight, since the classical SB problem assumes that particles dynamics are unaffected by external potential functions $U$. This assumption, however, limits our ability to model complex real-world physical systems.
Intermediate observations represented by marginal distributions $\{\rho_{m}\}_{m=1}^{M}$ at specific time steps $\{t_{m}\}_{m=1}^{M}$, can also be incorporated as additional constraints (chen2023deep; tang2025branched). This leads to the mmGSB problem,
@@
Wasserstein Hamiltonian flows and geodesics.   A convenient solution to the mmGSB problem (2) is to specify analytical optimality conditions (Sec. 2). These take the form of a Wasserstein Hamiltonian Flow (chow2020wasserstein), which describes a probability distribution evolving according to Hamiltonian dynamics. This evolution lies on planes tangent to $\mathcal{P}^{+}(\mathcal{M})$, specifically on the cotangent bundle $\mathcal{T}^{*}\mathcal{P}^{+}(\mathcal{M})$, the dual of $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$, and it is governed by the derivatives of a scalar Hamiltonian function $H$. However, their integration remains computationally expensive (buzun2025hotahamiltonianframeworkoptimal; wu2025parameterized), so we propose a geometric reformulation that results in a significant simplification. To this end, we introduce Proposition 1, a standard result from differential geometry (App. B.2), which is instrumental in lifting these equations to geodesics on $\mathcal{P}^{+}(\mathcal{M})$.
@@
Let the optimality conditions of the mmGSB problem (2) be expressed in Hamiltonian form, yielding the optimal bridge $\rho^{t}(x)$. Then, $\rho^{t}(x)$ can be viewed as a geodesic connecting the marginals in equation 2c w.r.t. the modified Riemannian metric $g_{J}$, known as the Jacobi metric (abraham2008foundations).
To access the Jacobi metric and determine the corresponding geodesic, we first derive the Hamiltonian optimality conditions of the mmGSB problem (2) using Lagrange multipliers (cui2024stochastic). This introduces a potential function $S^{t}(x)$, whose gradient defines the drift via ${v^{t}(x)}=\nabla_{x}S^{t}(x)$. The potential enforces the dynamic constraint (2b) within the cost functional (2a), whose first variation yields the Hamiltonian optimality conditions,
@@
with the corresponding Hamiltonian, $H({\rho}^{t},{S}^{t})=\mathcal{K}({\rho}^{t},{S}^{t})+\mathcal{F}({\rho}^{t})$, defined as the sum of a kinetic energy $\mathcal{K}$ and a potential energy $\mathcal{F}=-U-I$, dependent on the potential function $U$ and the Fisher information $I$. A detailed derivation of these dynamics and the full Hamiltonian function is in App. C.1. To handle boundary conditions (2c), the Hamiltonian dynamics (3) are typically integrated backward in time, where the solution at each intermediate point $(\rho^{t_{m}},S^{t_{m}})$ serves as the initial condition for the next segment (theodoropoulos2025momentum). The potential function ${S}^{t}(x)$ in equation (3a) is linked to the infinitesimal density variation $\partial_{t}{\rho}^{t}\in\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ via the weighted Laplacian operator $\Delta_{{\rho}^{t}}$, introduced through the Wasserstein metric (1). This connection establishes a correspondence between the tangent bundle $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ and the cotangent bundle $\mathcal{T}^{*}\mathcal{P}^{+}(\mathcal{M})$, where ${S}^{t}(x)$ naturally resides, and where the Hamiltonian dynamics of $({\rho}^{t},{S}^{t})$ unfold.

### v3

@@ 1 Introduction
Inferring the stochastic process that most likely generates a set of sparse observations is a fundamental challenge, e.g., in cellular dynamics (Yeo et al., 2021; Zhang et al., 2024; Moon et al., 2019), meteorology (Franzke et al., 2015), and economics (Kazakevičius et al., 2021; Huang et al., 2024). Here, the target is not merely the distributions of observed data, but rather the underlying dynamics of cell populations, weather patterns, or economic phenomena, enabling reconstruction of missing intermediate states and predicting the systems’ future evolution.
The Schrödinger Bridge (SB, Schrödinger (1931)) is a powerful mathematical framework to address this. SB seeks the most likely stochastic path between marginals (i.e., observations), while being close to a reference process, typically Brownian motion. This offers a general stochastic optimal control perspective that encompasses both Optimal Transport (OT, Vargas et al. (2021)) and generative approaches such as diffusion models (Ho et al., 2020; Chen et al., 2024), which can be interpreted as optimal bridges with a Gaussian initial marginal (Bortoli et al., 2021). Unfortunately, current SB solvers operate on an infinite probability space and rely on iterative forward–backward stochastic simulations or progressive refinement of the reference dynamics. This leads to complex and costly optimizations, limiting adoption.
@@
This paper overcomes the energy-preserving limitation and the iterative optimization schemes of existing SB approaches by formulating a novel geometric generalization of the SB to model non-conservative systems and by proposing a near-linear time solver. Specifically, we build on the geometric perspective of the SB, which casts it as a flow governed by Hamiltonian dynamics on the Wasserstein probability space (Sec. 3). Extending the Hamiltonian system to a contact Hamiltonian (Zadra, 2023), we propose a more general energy-varying formulation of the SB problem: The Non-Conservative Generalized SB (NCGSB, Sec. 4). To make computations tractable, we introduce the Contact Wasserstein Geodesic (CWG), which solves the NCGSB problem by casting it as a geodesic computation, reshaping its cost functional into a Riemannian metric whose induced distance is minimized. Discretizing the geodesic leads to geodesic segments that match standard residual blocks. We show how this leads to an efficient solver that avoids outer iteration loops and achieves near-linear complexity in both dimensionality and batch size. Additionally, our approach allows for guided generation, steering the learned process via a task-dependent loss, which in our geometric framework corresponds to modulating the Riemannian metric (Sec. 5). We demonstrate our approach on benchmarks and tasks such as LiDAR manifold navigation, molecular dynamics predictions, and image-based reconstruction of systems such as sea-surface temperature and robotic pick-and-place (Sec. 6).
In summary, we contribute: (1) A novel non-conservative formulation of the Schrödinger Bridge problem that models a wider range of real-world physical stochastic processes by building on geometric flows governed by contact Hamiltonian dynamics; (2) The introduction of the Contact Wasserstein Geodesic (CWG) framework, a general geometric solver compatible with all Schrödinger Bridge variants, enabling fast and scalable computation; (3) a guided generation method for steering the bridge learned by CWG according to new task specifications.
@@ 2 Related Work
Schrödinger bridges.   The SB problem imposes no constraints on the probabilistic path beyond matching the endpoint marginals, limiting its applicability when intermediate observations exist. It also does not permit including known physical laws governing the system’s dynamics. The multi-marginal Schrödinger Bridge (mmSB) treats intermediate observations (Theodoropoulos et al., 2025) as constraints, which enables reconstruction of continuous dynamics without piecewise approximations. In contrast, the Generalized Schrödinger Bridge (GSB) adds a state cost, allowing potential energy functionals to be minimized along the probability path. This let us to model mean-field interactions (Gaitonde et al., 2021; Ruthotto et al., 2020), conservative forces (Philippidis et al., 1979; Noé et al., 2020), or geometric priors (Chen and Lipman, 2024; Liu et al., 2018).
@@
Non-conservative Schrödinger bridge formulations.   The SB problem assumes constant energy, preventing it from modeling non-conservative systems. The momentum SB augments the state space with velocity (Theodoropoulos et al., 2025; Chen et al., 2023), allowing damping to be incorporated (Blessing et al., 2025; Sterling et al., 2025). This, however, doubles the state space and increases computational cost. Other extensions replace the Brownian reference process with the Ornstein–Uhlenbeck (OU) process (Orland, 2025; Zhang and Stumpf, 2025), introducing a non-conservative prior for the target dynamics. The OU process defines a mean distribution with a curl-free component that drives convergence and a divergence-free component that induces rotation. Yet, this formulation lacks a mechanism for energy dissipation in the rotational dynamics. We introduce a more general energy-varying framework in which dissipation naturally emerges across all components of the dynamics, while only requiring a scalar augmentation of the state space.
@@
Schrödinger bridge solvers.   Matching-based iterative approaches (Bortoli et al., 2024; Shi et al., 2023; Gushchin et al., 2024) have gained popularity by improving the scalability and robustness of traditional Iterative Proportional Fitting (IPF) algorithms (Kullback, 1968; léonard2013surveyschrodingerproblemconnections) through Markovian projections, thereby avoiding the need for full trajectory storage. However, for GSB, restrictive assumptions like Gaussian probability paths limit expressivity (Liu et al., 2024; Tang et al., 2025), while for mmSB, global trajectory consistency is compromised due to the piecewise nature of the approach and its sensitivity to the initial choice of the reference process (Shen et al., 2025). To overcome these limitations, the stochastic dynamics can be learned indirectly by leveraging the analytical optimality conditions of the SB problem (Vargas et al., 2021; Chen et al., 2022). This has been shown to scale more effectively to both the GSB (Liu et al., 2022; Buzun et al., 2025) and the mmSB problem (Theodoropoulos et al., 2025; Hong et al., 2025; Chen et al., 2023), due to the symmetries inherent in these optimality conditions (App. A). However, these methods are limited by a classical SB solver bottleneck: the computational overhead of their iterative nature, which alternates between forward–backward passes or repeated dynamics integration and reference updates. We propose a cheaper non-iterative solver that scales nearly linearly with both dimensionality and sample size.
@@
Schrödinger bridge guided generation.   The SB framework naturally extends to conditional generation, where the marginals and the transition path depend on additional parameters or objectives (Shi et al., 2022). Model guidance techniques (Song et al., 2023; Guo et al., 2024) introduce a guidance term into the stochastic process, typically derived from the gradient of a loss function. This approach steers the flow locally and resembles a gradient-based form of optimal control. Alternatively, Raja et al. (2025) employs a global optimal control perspective. Their approach generates full trajectories and chooses the one minimizing a task-specific action functional. However, their method is deterministic and yields a single optimal path rather than a posterior distribution over paths. Unlike previous guidance methods, we propose a hybrid approach to guided generation within the NCGSB framework. By embedding a task-specific loss into the potential, we reshape the Riemannian metric so that the resulting geodesic reflects the guidance objective, allowing the learned dynamics to align with the desired outcome.
@@ 3 Preliminaries
We introduce the necessary notation throughout the following sections, and summarize it in App. B.
The Wasserstein manifold.   Before formally introducing the SB problem, we define its domain. Let $\mathcal{P}^{+}(\mathcal{M})$ denote the space of smooth, positive density functions supported on a manifold $\mathcal{M}$. Each element $\rho\in\mathcal{P}^{+}(\mathcal{M})$ is a function $\rho(x):\mathcal{M}\to\mathbb{R}^{+}$ satisfying $\int_{\mathcal{M}}\rho(x)\,dx=1$. The density dynamics is represented by a time-dependent family $\{\rho^{t}\}_{t\in\mathbb{R}^{+}}\subset\mathcal{P}^{+}(\mathcal{M})$. The infinitesimal variation of the density at time $t$ is the time derivative $\partial_{t}\rho^{t}(x)$, which lies in the tangent space $\mathcal{T}_{\rho^{t}}\mathcal{P}^{+}(\mathcal{M})$. The collection of all such tangent spaces forms the tangent bundle $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$. When equipped with the Wasserstein metric, $\mathcal{P}^{+}(\mathcal{M})$ becomes a Riemannian manifold (Ambrosio et al., 2005). The corresponding metric tensor is defined as,
@@
where $(-\Delta_{\rho^{t}})^{\dagger}$ is the inverse of the weighted Laplacian operator $\Delta_{\rho^{t}}=-\nabla_{x}\cdot(\rho^{t}\,\nabla_{x})$ (Chow et al., 2020), inducing an inner product on $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ and a distance $d_{\mathcal{W}_{2}}(\rho_{a},\rho_{b})$ for $\rho_{a},\rho_{b}\in\mathcal{P}^{+}(\mathcal{M})$. The minimum-length curve $\rho^{t}$ connecting the two distributions $\rho_{a},\rho_{b}$ is called a geodesic.
Multi-marginal generalized Schrödinger bridge (mmGSB).   Given two endpoint densities $\rho_{a},\rho_{b}\in\mathcal{P}^{+}(\mathcal{M})$, the SB problem (Schrödinger, 1931; 1932) seeks the most probable interpolating density path $\rho^{t}$. This minimizes the Kullback–Leibler divergence w.r.t. a reference process $\rho^{t}_{\text{ref}}$, typically Brownian motion. The SB problem is equivalent to a stochastic optimal control setting (Dai Pra, 1991), which minimizes the cost required to transport a set of diffusing particles from an initial distribution $\rho_{a}$ to a target distribution $\rho_{b}$. This dynamic reformulation of the OT problem (Benamou and Brenier, 2000; Chen et al., 2014) has solutions corresponding to geodesics on the Wasserstein manifold $\mathcal{P}^{+}(\mathcal{M})$. These trajectories are straight, since the classical SB problem assumes that particles dynamics are unaffected by external potential functions $U$. This assumption, however, limits our ability to model complex real-world physical systems.
Intermediate observations represented by marginal distributions $\{\rho_{m}\}_{m=1}^{M}$ at specific time steps $\{t_{m}\}_{m=1}^{M}$, can also be incorporated as additional constraints (Chen et al., 2023; Tang et al., 2025). This leads to the mmGSB problem,
@@
Wasserstein Hamiltonian flows and geodesics.   A convenient solution to the mmGSB problem (2) is to specify analytical optimality conditions (Sec. 2). These take the form of a Wasserstein Hamiltonian Flow (Chow et al., 2020), which describes a probability distribution evolving according to Hamiltonian dynamics. This evolution lies on planes tangent to $\mathcal{P}^{+}(\mathcal{M})$, specifically on the cotangent bundle $\mathcal{T}^{*}\mathcal{P}^{+}(\mathcal{M})$, the dual of $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$, and it is governed by the derivatives of a scalar Hamiltonian function $H$. However, their integration remains computationally expensive (Buzun et al., 2025; Wu et al., 2025), so we propose a geometric reformulation that results in a significant simplification. To this end, we introduce Proposition 1, a standard result from differential geometry (App. C.2), which is instrumental in lifting these equations to geodesics on $\mathcal{P}^{+}(\mathcal{M})$.
@@
Let the optimality conditions of the mmGSB problem (2) be expressed in Hamiltonian form, yielding the optimal bridge $\rho^{t}(x)$. Then, $\rho^{t}(x)$ can be viewed as a geodesic connecting the marginals in equation 2c w.r.t. the modified Riemannian metric $g_{J}$, known as the Jacobi metric (Abraham and Marsden, 2008).
To access the Jacobi metric and determine the corresponding geodesic, we first derive the Hamiltonian optimality conditions of the mmGSB problem (2) using Lagrange multipliers (Cui et al., 2024). This introduces a potential function $S^{t}(x)$, whose gradient defines the drift via ${v^{t}(x)}=\nabla_{x}S^{t}(x)$. The potential enforces the dynamic constraint (2b) within the cost functional (2a), whose first variation yields the Hamiltonian optimality conditions,
@@
with the corresponding Hamiltonian, $H({\rho}^{t},{S}^{t})=\mathcal{K}({\rho}^{t},{S}^{t})+\mathcal{F}({\rho}^{t})$, defined as the sum of a kinetic energy $\mathcal{K}$ and a potential energy $\mathcal{F}=-U-I$, dependent on the potential function $U$ and the Fisher information $I$. A detailed derivation of these dynamics and the full Hamiltonian function is in App. D.1. To handle boundary conditions (2c), the Hamiltonian dynamics (3) are typically integrated backward in time, where the solution at each intermediate point $(\rho^{t_{m}},S^{t_{m}})$ serves as the initial condition for the next segment (Theodoropoulos et al., 2025). The potential function ${S}^{t}(x)$ in equation (3a) is linked to the infinitesimal density variation $\partial_{t}{\rho}^{t}\in\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ via the weighted Laplacian operator $\Delta_{{\rho}^{t}}$, introduced through the Wasserstein metric (1). This connection establishes a correspondence between the tangent bundle $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ and the cotangent bundle $\mathcal{T}^{*}\mathcal{P}^{+}(\mathcal{M})$, where ${S}^{t}(x)$ naturally resides, and where the Hamiltonian dynamics of $({\rho}^{t},{S}^{t})$ unfold.